Total population=60 million.

• Total labour force 40 million.

- Number of employed =30 million.

• Natural rate of unemployment 12%.

a. Find the total unemployment rate.

b. Calculate the cyclical unemployment rate.
1
Expert's answer
2021-03-05T11:17:30-0500

Unemployment rate = Number of Unemployed Persons / Labor Force

a) Unemployment rate = (40-30)/40 = 25%

Cyclical unemployment rate = Total unemployment rate - Natural rate of unemployment

b) Cyclical unemployment rate = 25-12= 13%